User Tools

Site Tools


lbaops:lbajul2011:v252zpalog

Parkes recording to ATNF V001B

Fringes just before experiment start, but oddly weak. Absent at first, had to change recorder to 3,4,7,8 mode to find them. Worried we might still have wrong polarisation?

Confirmed, polarisations swapped until 18:25. Corrected, fringes to Pa-Ho have the same strength as At-Ho now.

Rx change from 22:26–04:56 UT Off source and stopped recorder

lbaops/lbajul2011/v252zpalog.txt · Last modified: 2015/12/18 16:38 (external edit)